OpenStudy (anonymous):

The smallest package weighed only 1/3 more than 1/2 the weight of the heaviest package. If the two packages combined weighed 100 kilograms, how much did the smaller package weigh?

OpenStudy (anonymous):

let the heaviest package weigh = x kg 1/2 its weight = x/2 kg then the smallest package weighs = x/2 + 1/3*x/2 = x/2 + x/6 ATQ x/2 + x/2 + x/6 = 100 3x + 3x + x ------------- = 100 6 10x = 600 x = 600/10 = 60 so the heaviest package weighs = 60 kg smaller package weighs = 60/2 + 60/6 = 30 + 10 = 40 kg

OpenStudy (anonymous):

if satisfied with the ans, pls click GOOD ANSWER button o( :})